# Fourier PINNs: From Strong Boundary Conditions to Adaptive Fourier Bases

### v1

@@ 6.5 Discussion
We can see that in all examples for both the 1D and 2D problems, the standard PINN, W-PINN, and A-PINN failed to obtain reasonable solutions, with large relative $L_{2}$ errors for all cases (e.g.,$\sim 1.0$). This implies that these methods are unable to capture high-frequency signals. Additionally, RFF-PINN can achieve small relative errors with particular scale settings; however, for most settings (60-70% of the 20 total RFF-PINN settings listed in Table 1), RFF-PINN failed with large solution errors. These results show that the success of RFF-PINN is very sensitive to the choice of number and scales, and there is no apparent correspondence between scale settings and problem solutions. For example, for the single-scale true solution (see Figure 4), the successful cases of RFF-PINN include two, three or even five scale settings, e.g.,(1, 20, 194), yet it failed with all the single-scale settings. The spectral method also failed to obtain good solutions, implying that only estimating a linear combination of Fourier bases are not sufficient. As a comparison, Fourier PINNs achieves reasonable solution accuracy in all the cases, with relative $L_{2}$ errors between $10^{-3}$ and $10^{-4}$. Our method is also robust to the choice of the frequency range $K$. As long as $K$ is large enough, our method nearly always selected the target frequency, and returned similar solution accuracy. Note that, to focus on the evaluation of the key idea, Fourier PINNs did not employ any loss term re-weighting schemes to further improve the accuracy, such as the NTK re-weighting used in RFF-PINN or mini-max updates. However, it would be straightforward to integrate these methods into the proposed algorithmic framework.
Figure 14: 2D Poisson equation with true solution $u(x,y)=\sin(6x)\cos(20x)+\sin(6y)\cos(20y)$
Figure 15: Predicted solutions (left) and absolute errors (right) of each method on the 2D Poisson equation with true solution $u(x,y)=\sin(6x)\cos(20x)+\sin(6y)\cos(20y)$
@@
We finally visualize the solution of each method in solving a 2D Poisson equation (see Figure 12) to visualize the point-wise errors for each method within the two dimensional domain. The ground-truth solution is given in Figure 14 and the solution obtained by each method is shown in Figure 15. Note that for W-PINN and RFF-PINN, we report the result with the smallest relative $L_{2}$ error. While the spectral method also obtains good results, the point wise error biases much on the regions that are close to the boundary and the relative $L_{2}$ error is two orders of magnitude larger than Fourier PINN’s. RFF-PINN (with the best number and scale choice) can roughy capture the solution structure, but the accuracy is far worse than Fourier PINNs and the spectral method. For example, in Figure 12, RFF-PINN nearly failed with every choice of the number and scale set. We can see that overall, Fourier PINNs best recovers the original solution.
@@ 7 Conclusion
We have presented Fourier PINNs, a novel PINN extension to capture high-frequency and multi-scale solution information. Our adaptive basis learning and selection algorithm can automatically identify target frequencies and truncate useless ones, without the need for tuning the number and scales. Currently, we assume the spacing (granularity) of the candidate frequencies is fine-grained enough, which can be overly optimistic. In the future, we plan to develop adaptive spacing approaches to capture this granularity.
@@ Acknowledgments
We acknowledge the use of large language models (LLMs) for manuscript editing.

### v2

@@ 6.7 Discussion
The 1D and 2D problems results reveal that standard PINN, W-PINN, and A-PINN consistently fail to achieve reasonable solutions, with relative $\ell_{2}$ errors remaining large ($\sim 1.0$) across all cases. This indicates that these methods struggle to capture high-frequency signals effectively. While RFF-PINNs achieve lower relative errors under specific scale configurations, they exhibit significant sensitivity to the choice of the number and range of scales. For most configurations (60–70% of the 20 tested settings listed in Table 2), RFF-PINNs failed with large solution errors. Notably, even in the single-scale solution case (Figure 15(a)), successful configurations often involved multi-scale settings (e.g., (1, 20, 194)), whereas all single-scale settings failed. These findings highlight the lack of a clear relationship between scale configurations and solution accuracy, posing a significant challenge for RFF-PINNs. In contrast, Fourier PINNs consistently delivered accurate results, with relative $\ell_{2}$ errors ranging from $10^{-3}$ to $10^{-4}$ across almost all test cases. A key advantage of Fourier PINNs is their robustness to the choice of the frequency range $K$. Provided $K$ is sufficiently large, the method reliably selects the target frequencies, achieving comparable solution accuracy across cases. It is important to note that Fourier PINNs do not employ advanced loss re-weighting schemes (e.g., NTK re-weighting used in RFF-PINNs or mini-max updates) in this study, as the focus was on evaluating the core algorithm. However, such enhancements could be seamlessly integrated into the Fourier PINN framework to further improve accuracy. Additionally, Fourier PINNs outperform baseline methods in higher-dimensional problems, demonstrating scalability and the ability to handle high-frequency and multi-scale solutions. However, their performance on (some) nonlinear, higher-dimensional PDEs, such as the 2D Allen-Cahn example in Equation 41, highlights areas for future work. Specifically, improving training efficiency and refining the network architecture to better address the challenges posed by nonlinear dynamics and stiffness could further enhance the applicability of Fourier PINNs.
@@
We finally visualize the solution of each method in solving a 2D Poisson equation (see Figure 17(b)) to visualize the point-wise errors for each method within the two-dimensional domain. The ground-truth solution is given in Figure 20, and the solution obtained by each method is shown in Figure 21. While the spectral method also obtains good results, the point-wise error biases much on the regions close to the boundary, and the relative $\ell_{2}$ error is two orders of magnitude larger than Fourier PINN’s. RFF-PINN (with the best number and scale choice) can roughly capture the solution structure, but the accuracy is far worse than that of Fourier PINNs and the spectral method. For example, in Figure 17(b), RFF-PINN nearly failed with every choice of the number and scale set. We can see that Fourier PINNs best recover the original solution.
7 Future Work: Improving Scalability with Tensor Decomposition Techniques
One limitation of our current approach is its scalability in higher-dimensional settings. The computational cost grows exponentially with the problem’s dimensionality, making it infeasible for very high-dimensional cases. Future work could explore tensor decomposition techniques, such as Tensor-Train (TT) decomposition and Tucker decomposition, to reduce the computational complexity and memory requirements.
Specifically, in high-dimensional spaces, the solution $u(x_{1},x_{2},\ldots,x_{d})$ can be represented as a tensor-product over basis functions $\phi(x_{i})$, with the weights forming a high-dimensional tensor $\mathcal{U}$:
$u(x_{1},x_{2},\ldots,x_{d})\approx\sum_{i_{1}=1}^{n_{1}}\sum_{i_{2}=1}^{n_{2}}\cdots\sum_{i_{d}=1}^{n_{d}}\mathcal{U}_{i_{1},i_{2},\ldots,i_{d}}\,\phi_{i_{1}}(x_{1})\,\phi_{i_{2}}(x_{2})\cdots\phi_{i_{d}}(x_{d}).$ (44)
Without decomposition, storing and computing with $\mathcal{U}$ requires $\prod_{j=1}^{d}n_{j}$ parameters, which is impractical in high dimensions.
One promising approach to mitigate this is through Tucker decomposition, where $\mathcal{U}$ is decomposed into a core tensor $\mathcal{C}$ and a set of factor matrices $\{\mathbf{A}_{j}\}$:
$\mathcal{U}_{i_{1},i_{2},\ldots,i_{d}}\approx\sum_{j_{1}=1}^{r_{1}}\sum_{j_{2}=1}^{r_{2}}\cdots\sum_{j_{d}=1}^{r_{d}}\mathcal{C}_{j_{1},j_{2},\ldots,j_{d}}\,\mathbf{A}_{1}(i_{1},j_{1})\,\mathbf{A}_{2}(i_{2},j_{2})\cdots\mathbf{A}_{d}(i_{d},j_{d}),$ (45)
where each factor matrix $\mathbf{A}_{j}\in\mathbb{R}^{n_{j}\times r_{j}}$ maps the original tensor to a lower-dimensional representation. This decomposition reduces the number of parameters to $\sum_{j=1}^{d}n_{j}r_{j}+\prod_{j=1}^{d}r_{j}$, allowing scalability as it grows linearly with $d$ under fixed ranks $r_{j}$. Another promising approach is Tensor-Train (TT) decomposition, where $\mathcal{U}$ is expressed as a chain of lower-order tensors, known as TT-cores:
$\mathcal{U}_{i_{1},i_{2},\ldots,i_{d}}=\sum_{\alpha_{1}=1}^{r_{1}}\sum_{\alpha_{2}=1}^{r_{2}}\cdots\sum_{\alpha_{d-1}=1}^{r_{d-1}}\mathbf{G}_{1}(i_{1},\alpha_{1})\,\mathbf{G}_{2}(\alpha_{1},i_{2},\alpha_{2})\cdots\mathbf{G}_{d}(\alpha_{d-1},i_{d}),$ (46)
where $\mathbf{G}_{1}\in\mathbb{R}^{n_{1}\times r_{1}}$, $\mathbf{G}_{d}\in\mathbb{R}^{r_{d-1}\times n_{d}}$, and intermediate cores $\mathbf{G}_{j}\in\mathbb{R}^{r_{j-1}\times n_{j}\times r_{j}}$ for $j=2,\ldots,d-1$. The TT-ranks $r_{j}$ control the complexity of the representation, enabling a linear scaling of parameters with $d$ for fixed ranks.
@@ 8 Conclusion
We have presented Fourier PINNs, a novel extension to capture high-frequency and multi-scale solution information. Our adaptive basis learning and selection algorithm can automatically identify target frequencies and truncate useless ones without tuning the number and scales. We assume the spacing (granularity) of the candidate frequencies is fine-grained enough, which can be overly optimistic. In the future, we plan to develop adaptive spacing approaches to capture this granularity. Applying tensor decomposition techniques, such as Tucker and Tensor-Train, to the weight tensor $\mathcal{U}$ over basis functions could provide a viable path for extending our approach to higher dimensions. By reducing the effective dimensionality of the parameter space, these methods offer a promising direction to overcome scalability challenges and maintain the advantages of meshless approaches in high-dimensional spaces.
@@ Acknowledgments
We acknowledge the use of large language models (LLMs) for manuscript editing. This work has been supported by MURI AFOSR grant FA9550-20-1-0358, NSF CAREER Award IIS-2046295, and NSF OAC-2311685 and the Army Research Office (ARO).